find dy/dx when y= "(1+x\/1-x)" -2


1
Expert's answer
2022-02-07T14:21:10-0500

"\displaystyle
y=\left(\frac{1+x}{1-x}\right)^2"

Let "\displaystyle
p=\frac{1+x}{1-x}," then "\displaystyle
y=p^2"

From "\displaystyle
y=p^2,\ \frac{dy}{dp}=2p".

Also, from "\displaystyle
p=\frac{1+x}{1-x},\ \frac{dp}{dx}=\frac{(1-x)(1)-(1+x)(-1)}{(1-x)^2}=\frac{2}{(1-x)^2}", by quotient rule.

Thus by chain rule, we have that;

"\displaystyle
\frac{dy}{dx}=\frac{dy}{dp}\cdot\frac{dp}{dx}=2p\cdot\frac{2}{(1-x)^2}=2\left(\frac{1+x}{1-x}\right)\cdot\frac{2}{(1-x)^2}=\frac{4(1+x)}{(1-x)^3}"